OpenStudy (anonymous):

Morristown has a population of 18000. its population is decreasing at a rate of 1.2%. what will be the population at half if the p(t)=18,000 *0.988t

OpenStudy (tkhunny):

Do you mean WHEN will the population be 1/2 what it was? \[18000*0.988^{t} = 9000\] You might need a logarithm for htat.

OpenStudy (anonymous):

I dont have one of those, so at 0 will it be 18,000

OpenStudy (tkhunny):

Yes, you should have p(0) = 18000 and your task is to find a value, t, such that p(t) = ½(18000)

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

graphing, f(x)=-2*2x, how do I get started

OpenStudy (anonymous):

can someone tell me ho to graph f (x) =1/3x

OpenStudy (tkhunny):

Do you have a calculator wiht a "Table" feature? Or just the ability to draw graphs. It would be a worthwhile adventure through the owner's manual.